What would happen to the average kinetic energy of the molecules of a gas sample if the
temperature of the sample increased from 20°C to 40°C?
A) It would become half its value.
B) It would double.
C) It would decrease.
D) It would increase.
E) Two of these.
Answer #1

we know that

average kinetic energy of a molecule at a temperature is given by

K.Eavg = 1.5 x k x T

here

k is the boltzmann constant

T is the temperature in kelvin scale

so

at two differenet temperatures we get

K.E2/ K.E1 = T2 / T1

given

T1 = 20 C

T1 = 20 + 273 K

T1 = 293 K

also

T2 = 40 C

T2 = 40 + 273

T2 = 313 K

so

using those values

we get


K.E2/ K.E1 = T2 / T1

K.E2/ K.E1 = 313/293

K.E2/ K.E1 = 1.068


so

the average kinetic energy increased by 1.068 times


so the answer is

option D) it would increase